From GlueXWiki
Jump to: navigation, search

June 19, 2014 FDC+CDC meeting


  1. Gas system leakage test, [1], [2], [3], [4]
  2. DAQ (Beni)
  3. CDC status (Mike)
  4. Electronics (Chris, Nick)
  5. Readiness review (Beni, Lubomir)
  6. Engineering (Bill)
    • cathode measurements
  7. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Curtis, Naomi (CMU), Dave, Eugene, Chris, Nick, Simon, Fernando, and Lubomir (JLab).


- Nick: three CDC HV modules replaced with the ones recently repaired and they are working fine. In the moment we have 2 spare 1550P modules and another 2 are being repaired in Italy. In addition, Fernando ordered two extra 1550Ps that will be delivered by September.

- Two thermocouples have been installed on the CDC: one on the cooling duct and the other on the detector. Thus, one can check if the cooling is working. After some measurements it will be possible to interlock the CDC LV supply in case the cooling fails.

Gas system

- Leak test of the gas system has been performed by closing the mass flow controllers to the detectors first, then closing all the inputs and outputs of the two mixing tanks (pressure graph linked above). It showed a drop of ~1psi in four hours for the CDC tank and much less for the FDC tank. The sniffer showed a leakage on the CDC relief valve of ~10^-3 (from 10^-5 in air) and ~10^-4 on the FDC relief valve. Scott ordered relief valves with better quality from Circle Valve Technology Inc.

- We used this test to check the leaks in the FDC. The pressures in each cell of the four packages are linked above. The first three packages showed no measurable leak in 6 hours. For this time, the pressure in the fourth package decreased from 60 Pa to 36Pa in all the six cells. The leak is not significant and most likely it is outside of the detector (same for all the cells). We will check it once we have an access from the downstream side of the magnet.


- Beni has the DAQ on the CDC working: looking at the raw fADC data one certainly sees cosmic pulses. Beni is working now on the translation tables to be able to reconstruct cosmic tracks. Two FDC crates had problems with the TIs but have been fixed.

Readiness review

- We went through the readiness review documents: the procedures are ready for the CDC, FDC, and the gas system. The checklist is ready for the FDC only. Fernando: all the documents needed for the electronics are ready, as well. Fernando is ready also with the Electronics section of the Detector Design Report and will submit it also as a GlueX document.


- The plan for the cathode scan is to use the stringing table with the stage, motor and controller as originally used in Blue Crab. We will use the existing software just to move the stage with the microscope attached.

- Josh Foyles put the chiller on the scale again: 110-111 lbs fluctuating (calibration weight steady at 662 lbs). A month ago it was ithe same 411 lbs. Then the only way to explain the drop in the Fluorinert level is that it goes to another tank within the chiller. Bill will look into this.